36+7x-8y=0
-10y=-12-6x

Answer by mananth(16949)

7 x -8 y = -36 .............1
6 x -10 y = -12 .............2
Eliminate y
multiply (1)by -5
Multiply (2) by 4
-35 x 40 y = 180
24 x -40 y = -48
Add the two equations
-11 x = 132
/ -11
x = -12
plug value of x in (1)
7 x + -8 y = -36
-84 + -8 y = -36
-8 y = -36 + 84
-8 y = 48
y = -6